(a)
An angle measures 42. What is the measure of its supplement?
: (a)
An angle measures 42. What is the measure of its supplement?

Answer by jim_thompson5910(9926)

Let x=measure of supplement angle

Supplemental angles add up to 180 degrees. So that means that 42+x=180


42+x=180 Start with the given equation.


x=180-42 Subtract 42 from both sides.


x=138 Combine like terms on the right side.


----------------------------------------------------------------------

Answer:

So the answer is x=138 which means that the measure of the supplement angle is 138 degrees